Let {an} ∞ n=1 be a non-decreasing (resp. non-increasing) sequence which converges to a. Then prove that an ≤ a (resp. a ≤ an) for every n ∈ N.


1
Expert's answer
2021-02-02T04:37:20-0500

Suppose first "a_n" to be non-decreasing (the other case is totally symmetric). Suppose the contrary of the announced proposition, "\exists a_{N_0}, a_{N_0} >a". As "a_n" is non-decreasing "\forall n\geq N_0, a_n >0". Now by taking "\epsilon = |a-a_{N_0}|>0" we have by definition of convergence "N\in \mathbb{N}" such that "\forall n \geq N, |a_n-a|<\epsilon" . Now by taking "N'=\max (N_0, N)", we get "\forall n\geq N'" both properties : "a_n \geq a_{N_0} = a + \epsilon", "|a_n-a|<\epsilon". But this is clearly a contradiction, as the first inequality gives us "a_n-a \geq \epsilon". Therefore by contradiction "\forall n, a_n \leq a". The case of non-increasing sequence can be studied by considering the sequence "-a_n" that is non-decreasing and converges to "-a".